NAOJ GW Elog Logbook 3.2

Yuheng, Michael
We continued looking at the CC PLL bandwidth using the summing amplifier transfer function measurement.
I saw an old post by Marco Vardaro which says that the nominal bandwidth of the PLL is 40-50 kHz, though maybe the settings were much different back then. Analog devices datasheet says the loop bandwidth is about 20 kHz (fig 1). In this case it was just inspected from the width of the spectral peak, so not particularly accurate.
We took open loop transfer function by checking the coherence in FFT mode (fig 2). In this case, the frequency response is achieved by injecting white noise and taking Ch2/Ch1. We start with a small noise value and increase it until the coherence is approximately 1 over the frequency band of interest. For the CC OLTF the maximum noise excitation is about 30 mV. Also for whatever reason the spectrum analyzer initially gave me frequency response magnitude in dB/rtHz which was a bit odd, I don't know what setting was doing that. It went back to just dB after I switched to swept sine and then back. The unity gain frequency is about 11 kHz and the unity gain phase is +50 degrees (fig 3), which is a bit strange. Judging by this as well as the earlier PLL data it seems there is room to increase the bandwidth. We also checked in swept sine mode for which the optimal excitation amplitude was 15 mV. It gave almost the same result 11.2 kHz UGF.
Afterwards I noticed that this was actually measuring both the fast and slow loop. We repeated the measurement again with only the fast loop. In this case the optimal swept sine amplitude dropped to 10 mV and the UGF very slightly decreased to 10.8 kHz (fig 4, 5).
We used the ppol PLL as a reference check. Looking at the fast loop we see a unity gain frequency of 9.6 kHz with phase 50 degrees. In this case the optimal noise excitation was 150 mV and optimal swept sine was 100 mV (fig 6, 7). Even though the unity gain frequency is the same, it seems the ppol PLL is more robust against unlock. In this case the possible issue might not be control bandwidth but rather the dynamic range of actuation.
While searching for elogs about PLL bandwidth I came across a previous post by myself and Yuhang. At that time it seems like the correction signal for the CC PLL was too small. This was when we were having small glitch issues and not the major instability. In that case it seems that when the CC PLL was left floating 2 MHz off the setpoint, the fast correction signal was only 200 mV, versus the CC laser PZT tuning coefficient 1 V/MHz. At the time the ppol PLL fast correction signal gave the correct value. When I tried it today, CC detuning of about 7 MHz and ppol detuning of 3 MHz maxed out the fast correction signal. This is a bit strange so we should check again in detail tomorrow.

Aso, Yuheng, Michael
There is some indication that the CC PLL has a weak lock, so we tried to make an open loop transfer function measurement. Unlike the cavity servos however, the CC PLL servo doesn't have a perturb or monitor ports.
Takahashi-san gave us an old generic summing amplifier (from the original TAMA end mirror) which can be inserted into the CC PLL feedback path to inject swept sine or noise. It has IN, OUT, ADD which function like EPS1, EPS2, PERTURB IN on the cavity servos, and then monitor ports for the input and output signal. There are four sets of amplifiers named after IN/END PITCH YAW.
We checked the frequency response of the summing amplifier. With nothing in ADD, the IN MON, OUT MON and OUT signals should all be the same as IN. We tried first 3 Hz 1Vpp, which was what the function generator was set to. OUT and OUT MON could reproduce the signal but IN MON was attenuated. When the signal was a ramp, IN MON also had some integration, so the summing amplifier is not so useful for low frequency (fig 1, 2). But we don't really care so much about low frequency to check the PLL fast loop. We then tried 10 kHz 1Vpp which was reproduced in all outputs so it seems fine for higher frequency (fig 3). 60 kHz 0.2 Vpp was sent to ADD and it seems to have the proper response for the output and monitor channels (fig 4).
After testing we set up as follows. CC PLL FAST outputs the fast correction signal to the laser PZT. We connected this to IN, then connected OUT to the laser PZT. ADD, IN MON and OUTMON go to Source, Ch2, Ch1 on the SR785 spectrum analyzer (fig 5). The PLL was locked and then we checked the transfer function up to 100 kHz for swept sine. The excitation frequency that could be applied before unlock was quite small, about 10 mV. But we didn't see anything really meaningful. We checked in FFT mode using noise excitation but couldn't get meaningful coherence anywhere in the full span of the spectrum analyzer. I should more carefully check the relevant frequency range and actuation strength of the CC PLL fast signal. The SR785 only goes up to 100 kHz. The other spectrum analyzers we use to monitor the PLL can look at up to 2 GHz but are only single channel.
When we locked the PLL, I compared the time traces with what I saw previously to show the CC glitch problem (fig 6, from before). Only to see that the CC glitches did not appear (fig 7)... I don't have any idea what changed.

The calibration of new BS is shown in plot 1.
Input pol power has been normalised by the input laser power.
all other powers have been normalised by both the input laser power and the input power after the LC (which will be referred as input pol power).
Finally, I also normalise the input pol power by itself to have a unit 1 comparison for powers.

Aso-san showed me how to properly replace the potentiometers in the servo boards. So I hoped it was just the knob that was broken and not the variable resistor itself.
I took a replacement control knob from the elec shop, there's a lot of them.
Before reattaching I rotated the knob to about 5 out of 10, since the threshold was previously stuck at 0 volts, which is the middle of the range setting.
Unfortunately, the threshold out reading is still broken, as is the SHG lock. Threshold out reads zero normally, but when the dial is rotated there is a threshold voltage reading the magnitude of which depends on the rotation speed. So I think the potentiometer and threshold power supply connection work and there is a small circuit break somewhere common to the error signal and Threshold out reading, which is small enough such that the capacitance can detect a fast change in the potentiometer voltage divider.
I suspect the circuit break is at the marked location since neither the THRESHOLD OUT nor the SHG lock work correctly when the threshold dial is adjusted over the full range. However, THRESHOLD OUT can detect fast changes in the value, so the power supply connection and potentiometer probably still work.
Unfortunately the suspect point is the connection to screw terminal J14 and resistor R104 which is in a very inconvenient location.
SHG threshold and Auto Lock magically fixed itself and I don't know how or why. It even returned to the same value I had set (9 V). Output green is 310-320 mW, same as before.

Previously I had some issue with even basic shot noise spectrum being a bit glitchy, but actually it was just sensitivity to the ceiling lights. The LO shot noise spectrum is fine. Dark noise still has 50 Hz at odd numbered harmonics even with the homodyne plugged into the same power supply as the spectrum analyzer. Maybe I should try spectrum analyzer floating ground

I modified the circuit to match impedance with mokugo. Seems moku go only has impedance of 1M ohm and not 50ohm. So, I added 1Mohm (previously 50ohm) infront of Opamp and 10Mohm (previously 1.1k) on the feedback. I measured the voltage output after the opamp (CH1) and before the EOM (CH2). The EOM was connected to the circuit during this measurement.
The output doesn't match the expectation. Since I input 0.05V p-p from moku, I should get 0.5V after Opamp circuit (coz its gain is 10). But I get output as 8V both after the opamp circuit and before EOM. The picture is from the oscilloscope.
Maybe its because I left ground floating when I soldered the SMA cables.

We sometimes have a power ratio of more than 1 in our measurements. There could be several reasons for this, laser switched off between two measurements, some fluctuations after LC, etc. Hence, another BS (BS014) is added after LC to monitor some fluctuations that were not previously accounted for. I placed the power meter after 7 holes in the reflection of BS, and the camera after 15 holes from the transmission of BS. Sometimes, we require to know the size of the beam as well to take into account the power density.
The LC voltage change from 0-3.5V, with 0.1V step. Another power meter has been incorporated into the labview, and therefore 2 new columns has been added to the saving file. The last two column have the information from mean and std of power from this power meter.
1. After the BS with power meter in reflection and camera in transmission
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Polarization states\Fri, Nov 15, 2024 3-15-13 PM.txt
2. Only the power meter is after LC for this measurement, just to take into account the input power at each voltage. We only care about polarization states after the BS.
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Polarization states\Sat, Nov 16, 2024 3-40-15 PM.txt
3. Measurement with polarizer at two rotated angles (rotated about its optic axis)
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\GTPC Polarizer\20241105\45 deg\Fri, Nov 15, 2024 4-50-37 PM.txt
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\GTPC Polarizer\20241105\0 deg\Fri, Nov 15, 2024 6-24-29 PM.txt
4. Measurement with HWP at two rotated angles (rotated about its optic axis)
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\Retarder\HWP\20241115\0 deg\Fri, Nov 15, 2024 8-05-03 PM.txt
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\Retarder\HWP\20241115\38 deg\Fri, Nov 15, 2024 9-33-43 PM.txt

The SHG servo threshold setting is broken and outputs zero volts. This makes the SHG incapable of good green output, since it will just lock to zero volts output. Obviously no green means no squeezing so this needs to be fixed.
Turning the knob makes the signal jump a little with the polarity depending on the direction of adjustment. But then it goes back to zero.
Also the knob itself is broken and seems detached from the circuit assembly.
I guess the potentiometer connection is broken, but turning the knob maybe applies a little pressure on the circuit that is then released when I stop adjusting. Initially I thought the circuit was preferring to go through some capacitor but looking at the schematic it seems not likely.

I turned on all of the remaining filter cavity systems in the center and end room. Also Koach filter and air conditioning in ATC.
I reloaded the DDS. I didn't check everything yet, but the RF signal levels I did look at were the same as before.
Homodyne PSU: +19 V 64 mA, -19 V 60 mA ok
OPO Temp Act: 7.116 ok
SHG Temp Act: 3.191 ok
Clock 1: -10 dBm
Clock 2: -13 dBm
Clock 3: -11 dBm
Clock 4: -9.9 dBm
SHG EOM: -9.9 dBm (same as before, goes to +14 dB amplifier)
SHG demod: 4.5 dBm (same as before)
SHG mode matching: 1840 mV + (200, 140) -> 84.4% (same as before)
SHG demod phase: 110 -> 65
SHG error: 576 mVpp
SHG threshold: 5 V -> 9 V - had trouble locking to TEM00 at 5 V but at 9 V it was fine. As before, SHG threshold is 10x higher than the TRANSMISS OUT reading would suggest.
SHG green level: 330 mW (same as before)
GRMC was quite misaligned so I realigned.
GRMC mode matching: 1460 mV + (50, 50, 32, 24) -> 90.3%
GRMC demod phase: 135 -> 195
GRMC error: 1.68 V - the error signal is quite fuzzy. This is to be expected since the modulated sidebands from the ML EOM have to pass through three dichroics. But the error signal PD is very sensitive so the signal is large. I didn't see any of the strange behaviour of GRMC changing optimal demodulation phase at random, and I think it might have been me making an error in aligning the GRMZ assembly at one time. Maybe I forgot to turn on GRPS or MZ PZT while aligning it.
When I was just about to check the GRMC lock point, I lost the transmission, and the cause was another issue...

Measuring polarizer at 45 deg.
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\GTPC Polarizer\45 deg\Mon, Nov 11, 2024 8-22-25 PM.txt

the results (birefringence and diattenuation) for both points are attached.
Fig 1 , 2 - point where laser is at center
Fig 3, 4 - point where laser is on one side (not center)

I restarted wifi and ethernet, all hepa filters, old DGS and control pcs
The function generator used for timing is making a strange sound and should be monitored and probably replaced.

It seems I was supposed to put some voltage values. I didn't put any. Perhaps shouldn't have turned on. FYI I only turned on the power switch on the voltage extension board, which was outside for the FC one, and inside for the PCI one.

Workstation in on now.

1. Everything on bigfoot table
2. Server rack near pci
3. server rack near FC clean room
But, there is no internet(both wifi and ethernet) in tama right now. I don't see the ethernet on the pc. I don't know where to turn it on from.
It seems I was supposed to put some voltage values. I didn't put any. Perhaps shouldn't have turned on. FYI I only turned on the power switch on the voltage extension board, which was outside for the FC one, and inside for the PCI one.
I restarted wifi and ethernet, all hepa filters, old DGS and control pcs
The function generator used for timing is making a strange sound and should be monitored and probably replaced.

I wish to compare the effect of polcam setting on uncertainity (for future measurements), so I generated polarization states with different settings.
Polarization states are measured without BS
1. pol cam with full wave plate rotat 2048 pt fft, scan speed 60Hz
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Polarization states\20241106\Wed, Nov 6, 2024 8-32-38 PM.txt
2. pol cam with dual full rotation and 2048 pt fft, scan speed 60Hz
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Polarization states\20241106\Thu, Nov 7, 2024 8-05-47 AM.txt
3. full wave plate rotation by 1048pts, scane speed 60Hz (all the previous analysis were done using this setting)
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Polarization states\20241106\Fri, Nov 8, 2024 1-44-00 PM.txt
4. Polarizer was installed and measured using generated states
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\GTPC Polarizer\Fri, Nov 8, 2024 3-25-19 PM.txt
Fig 1 shows the orientation of the polarizer.
Measuring polarizer at 45 deg.
C:\Users\atama\Dropbox\LC-Experiment\Measurement Data\Birefringence Measurements\GTPC Polarizer\45 deg\Mon, Nov 11, 2024 8-22-25 PM.txt


In preparation for NAOJ blackout I completely shutdown all equipment in the ATC ISO Class 1 booth
Air conditioner, lightwave PSU, coherent PSU, SHG temperature controller, OPO temperature controller, homodyne PSU, PZT driver, Anritsu function generator, OPO translation driver, topgun
Then I disconnected all wall plugs for cleanroom equipment
Koach was already off

In preparation for NAOJ blackout, I completely shutdown all filter cavity related electronics.
All cleanroom equipment - lasers, NIM racks, photodetector power supplies, measurement instruments, computer, homodyne PSU, AOM RF amplifier PSU, rack RF amplifier PSU, filter fan. Then I switched off the power at the cluster behind PR/IMC_IN, which is the cleanroom main power.
External equipment - all oplev lasers, PD/QPD, coil drivers and pico controllers (PR, BS, IN, END), SR560, NIM racks, ADC, camera adapters, workstation, monitor, air conditioning main control

More detail later
I tried to find the source of the small squeezing instability, about 1 dB noise floor glitch every second or so. Initially I thought it was from ppol to OPO - I could see weird flashing that was quite frequent. When I was aligning ppol to the OPO I could also see higher order modes appear/disappear at about the glitch frequency. Also, I thought it would be related to ppol because the CC PLL and green controls seemed stable enough, and the homodyne spectrum shows no glitches when only IRMC is locked.
Actually, the ppol transmission spectrum flickering disappears when the SHG is unlocked. I also now notice the same behaviour of the glitchy transmission spectrum when observing OPO transmission of only CC and only main laser. So it's maybe some stray green problem.
I tried to look in squeezing again and somehow the squeezing spectrum didn't show the glitch issue anymore, however, depending on the spectrum analyzer span, there were a lot of noise peaks. So we still have some unclear noise issues.
Then it was time for electrical shutdown so I didn't have time to narrow down the problem further.